OpenStudy (loser66):

how to solve r^8+8r^4+16=0?

OpenStudy (mertsj):

(r^4+4)(r^4+4)=0

OpenStudy (loser66):

then?

OpenStudy (mertsj):

r^4=-4 r= fourth roots of -4

OpenStudy (calculusfunctions):

Yes now we will have complex roots. In other words, are roots will be imaginary. Have learned complex numbers? Have you been introduced to imaginary roots?

OpenStudy (loser66):

yes,

OpenStudy (calculusfunctions):

Alright then you know that the imaginary number\[i =\sqrt{-1}\]Correct?

OpenStudy (loser66):

yes

OpenStudy (loser66):

I need all of 8 solutions.

OpenStudy (calculusfunctions):

So then then in general, the sum of two squares may be written as\[a ^{2}+b ^{2}= (a +bi)(a -bi)\]

OpenStudy (calculusfunctions):

Hence you can factor each factor further by this principle.

OpenStudy (loser66):

I factor and it repeats many time while I need 8 separate solutions for my differential equations. :(.

OpenStudy (calculusfunctions):

Right so you had\[(r ^{4}+4)(r ^{4}+4)=0\] \[(r ^{2}+2i)(r ^{2}-2i)(r ^{2}+2i)(r ^{2}-2i)=0\]Do you understand thus far?

OpenStudy (loser66):

yes,

OpenStudy (calculusfunctions):

Now factor each of these factors and you will have 8 factors.

OpenStudy (loser66):

I will work on it, you still be here, right? I will show you my work and you will check it for me, right?

OpenStudy (calculusfunctions):

I'm here for now, I'll wait a while for you. I'll help someone else mean while. When you're ready, just ask for me.

OpenStudy (loser66):

thanks , but it is not the case of r^8 = -1, so I cannot apply e^(ipi) right?

OpenStudy (calculusfunctions):

right

OpenStudy (loser66):

thanks, I will tag you later.

OpenStudy (calculusfunctions):

OK

OpenStudy (loser66):

\[r^2+2i=0\\r^2=-2i\\r=\pm\sqrt{2i}\]and it repeats with r^2 -2i =0. HA!! I don't get.

OpenStudy (calculusfunctions):

For \[r ^{2}-2i =0\] \[r ^{2}=2i\] \[r =\pm \sqrt{2i}\]

OpenStudy (calculusfunctions):

For \[r ^{2}+2i =0\] \[r ^{2}=-2i\] \[r =\pm \sqrt{-2i}\]If you wish to simplify further,\[r =\pm i \sqrt{2i}\] \[r =\pm i ^{\frac{ 3 }{ 2 }}\sqrt{2}\]

OpenStudy (loser66):

your avatar looks exactly tcarroll010's one. At the first see, I tried to send message to you but cannot since you set your profile just receive message from person you fanned. I got mad with tcarroll010. I thought 'what is he doing?" hihihi.... when figure out the misunderstanding, I sent his message to let him know you "stole" his avatar.

OpenStudy (loser66):

but just 4 of them, not 8

OpenStudy (calculusfunctions):

I think he stole mine. Because I have had the same one for two years now.

OpenStudy (calculusfunctions):

Right, four equal and another four equal. For a total of eight.

OpenStudy (loser66):

I joined the site almost 1year, I didn't see you. Ha!!

OpenStudy (calculusfunctions):

Just like (x - 3)(x - 3) = 0 has two equal and real roots.

OpenStudy (calculusfunctions):

As a teacher, I'm busy, and thus not always available.

OpenStudy (calculusfunctions):

What grade are you in?

OpenStudy (loser66):

ok, got you. so, they are repeated and the solution for differential equations differentiated by multiply with polynomial. Thank you very much